OpenStudy (anonymous):

The perimeter of a rectangle is 54 cm. The area of the same rectangle is 176 cm². What are the dimensions of the rectangle?

OpenStudy (anonymous):

11 cm by 16 cm 8 cm by 22 cm 5.5 cm by 32 cm 4 cm by 44 cm

OpenStudy (anonymous):

2(l+b)=54 l*b=176 b=27-l l(27-l)=176 27l-l^2-176=0 l^2-27l+176=0 l^2-16l-11l-176=0 l(l-16)-11(l-16)=0 (l-11)(l-16)=0 l=11 or 16 b=16 or 11 so option A